OpenStudy (anonymous):

Which should you substitute as U when evaluating the integral of sin^4xcos^5xdx

OpenStudy (anonymous):

sinx=a

OpenStudy (anonymous):

a^4 . (1-a^2)^2. da is the integral.

OpenStudy (anonymous):

this is (a^6)/6 - (a^4)/4 + (a^2)/2 with a = sin x

OpenStudy (anonymous):

i.e U=sin x

OpenStudy (anonymous):

\[\sin ^4x (1-\sin^2x)^2cosxdx\]